Question 19
ProbabilityA bag contains 2 red and 4 green sweets of the same size and shape. Two boys pick a sweet each from the box, one after the other, without replacement. What is the probability that at least a sweet with a green wrapper is picked?
- A.1/5
- B.2/5
- C.8/15
- D.14/15
Show the answer and explanation
Answer: D (14/15)
Explanation
P(no green) = P(both red) = 2/6 × 1/5 = 1/15. So P(at least one green) = 1 − 1/15 = 14/15.